Modern pennies are composed of zinc coated with copper. A student determines the mass of a penny to be 2.484 g and then makes several scratches in the copper coating (to expose the underlying zinc). The student puts the scratched penny in hydrochloric acid, where the following reaction occurs between the zinc and the HCl (the copper remains undissolved):
Zn(s)+2HCl(aq)→H2(g)+ZnCl2(aq)
The student collects the hydrogen produced over water at 25 ∘C. The collected gas occupies a volume of 0.899 L at a total pressure of 792 mmHg .

Calculate the percent zinc in the penny. (Assume that all the Zn in the penny dissolves.)
Answer #1

Temperature, T = 273 + 25 = 298 K

Volume, V = 0.899 L

Pressure,P=total pressure - aqueous tension at 250C=792-23.8=768.2 mmHg=768.2/760=1.01atm

R = 0.0821 L.atm.K-1.mol-1

n = ?

Assuming ideal behaviour of H2 gas,

Ideal gas equation,

P V = n R T
n = 1.01 * 0.899 / (0.0821 * 298)

n = 0.0371 mol of H2 gas is produced.

From the balanced equation,

2 mol of H2 is produced from 1 mol of Zn

Then, 0.0371 mol of H2 is produced from 0.0371 * 1 / 2 = 0.01855 mol of Zn

Therefore, Mass of Zn = 0.01855 * 65.37 = 1.21 g.

% of Zn in penny = (1.21 / 2.484) * 100 = 48.7 %